A) Hydroxylamine hydrochloride is a powerful reducing agent which is used as a polymerization catalyst. It contains 5.80 mass % H, 20.16 mass % N, 23.02 mass % O, and 51.02 mass % Cl. What is its empirical formula? Determine the molecular formula of the compound with molar mass of 278 g.
Answer #1

1)

we have mass of each elements as:

H: 5.8 g

N: 20.16 g

O: 23.02 g

Cl: 51.02 g

Divide by molar mass to get number of moles of each:

H: 5.8/1.008 = 5.754

N: 20.16/14.01 = 1.439

O: 23.02/16.0 = 1.4387

Cl: 51.02/35.45 = 1.4392

Divide by smallest to get simplest whole number ratio:

H: 5.754/1.4387 = 4

N: 1.439/1.4387 = 1

O: 1.4387/1.4387 = 1

Cl: 1.4392/1.4387 = 1

So empirical formula is:H4NOCl

Answer: H4NOCl

2)

Molar mass of H4NOCl,

MM = 4*MM(H) + 1*MM(N) + 1*MM(O) + 1*MM(Cl)

= 4*1.008 + 1*14.01 + 1*16.0 + 1*35.45

= 69.492 g/mol

Now we have:

Molar mass = 278.0 g/mol

Empirical formula mass = 69.492 g/mol

Multiplying factor = molar mass / empirical formula mass

= 278.0/69.492

= 4

So molecular formula is:H16N4O4Cl4

Answer: H16N4O4Cl4